# Einstein from Noise: Statistical Analysis

### v2

@@ E.1 Auxiliary statements
$\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]-\frac{\sum_{r=0}^{d-1}f_{1}(r)e^{\tilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\tilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}0,$ (E.29)
@@
$\mathbb{E}\left[\left.{f_{2}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]-\frac{\sum_{r=0}^{d-1}f_{2}(r)e^{\tilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\tilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}0.$ (E.30)
Next, we evaluate the terms at the left-hand-side of (E.29) and (E.30).
Proof of (E.25) and (E.8).
We begin by proving that
$\frac{1}{2a_{d}}\frac{\sum_{r=0}^{d-1}{\tilde{\mu}_{k}[r]\exp\{\tilde{\mu}_{k}[r]a_{d}}\}}{\sum_{r=0}^{d-1}{\exp\{\tilde{\mu}_{k}[r]a_{d}}\}}\frac{\sigma_{k}^{2}}{\left|\mathsf{X}[k]\right|^{2}}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (E.31)
From the definition of $f_{1}(r)$, it follows that
$\sum_{r=0}^{d-1}{f_{1}(r)}=0,$ (E.32)
almost surely. Additionally, from the definition of $\tilde{\mu}_{k}[r]$,
${\tilde{\mu}_{k}[r]a_{d}}=2\sigma_{k}^{-1}a_{d}\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\cos\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right).$ (E.33)
By Assumption 4.2, we have $a_{d}\left|\mathsf{X}[k]\right|\to 0$ as $d\to\infty$. Thus, from (E.33) and the continuous mapping theorem, we obtain
${\tilde{\mu}_{k}[r]a_{d}}\xrightarrow[]{\mathsf{a.s.}}0.$ (E.34)
Since $\sum_{r=0}^{d-1}\tilde{\mu}_{k}[r]=0$ almost surely (from (E.32)), and applying the continuous mapping theorem along with (E.34), we deduce that
$\frac{{\sum_{r=0}^{d-1}{\tilde{\mu}_{k}[r]\exp\{\tilde{\mu}_{k}[r]a_{d}}\}}}{a_{d}\sum_{r=0}^{d-1}[{\tilde{\mu}_{k}[r]}]^{2}}\xrightarrow[]{\mathsf{a.s.}}1.$ (E.35)
Similarly, applying the continuous mapping theorem using (E.34), we get
$\frac{{\sum_{r=0}^{d-1}{\exp\{\tilde{\mu}_{k}[r]a_{d}}\}}}{d}\xrightarrow[]{\mathsf{a.s.}}1.$ (E.36)
Next, from the definition of $\tilde{\mu}_{k}[r]$,
$\sum_{r=0}^{d-1}[{\tilde{\mu}_{k}[r]}]^{2}=4\sigma_{k}^{-2}\left(\left|\mathsf{X}[k]\right|\left|\mathsf{N}[k]\right|\right)^{2}\sum_{r=0}^{d-1}\cos^{2}\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right).$ (E.37)
As $d\to\infty$,
$\frac{1}{d}\sum_{r=0}^{d-1}\cos^{2}\left({\frac{2\pi k}{d}r+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]}\right)\xrightarrow[]{\mathsf{a.s.}}\frac{1}{2}.$ (E.38)
Combining (E.37) and (E.38), we obtain
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}}\sum_{r=0}^{d-1}[{\tilde{\mu}_{k}[r]}]^{2}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2},$ (E.39)
for $d\to\infty$. Thus, combining (E.35)–(E.39), we conclude that
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}a_{d}}\frac{\sum_{r=0}^{d-1}\tilde{\mu}_{k}[r]e^{\tilde{\mu}_{k}[r]a_{d}}}{\sum_{r=0}^{d-1}e^{\tilde{\mu}_{k}[r]a_{d}}}\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (E.40)
This proves (E.31). Finally, combining (E.28), (E.29), and (E.40), we arrive at
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}a_{d}}\mathbb{E}\left[\left.{\tilde{\mu}_{k}[\hat{\mathsf{R}}]}\right|\mathsf{N}[k]\right]\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (E.41)
[…]
$\mathbb{E}\left[\frac{\sigma_{k}^{2}}{2a_{d}}\frac{\mathbb{E}[\mu(\hat{\mathsf{R}})|\mathsf{N}[k]]}{\left|\mathsf{X}[k]\right|^{2}}\right]\rightarrow\mathbb{E}\left[\left|\mathsf{N}[k]\right|^{2}\right].$ (E.42)
By combining (E.42) with (E.28), we get,
$\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}{\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\left|\mathsf{N}[k]\right|^{2},$ (E.43)
or equivalently,
$\mathbb{E}\left[\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]\right]\to\mathbb{E}\left[\left|\mathsf{N}[k]\right|^{2}\right],$ (E.44)
which proves (E.25).
By the law of total expectation, we then obtain
$\frac{\sigma_{k}}{a_{d}\cdot{\left|\mathsf{X}[k]\right|}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]\to\mathbb{E}\left[{\left|\mathsf{N}[k]\right|^{2}}\right]=2\sigma^{2},$ (E.45)
as $d\to\infty$. By Lemma E.2, we know that $\sigma_{k}^{2}\to 2\sigma^{2}$ as $d\to\infty$, so from (E.45), we conclude that
$\frac{1}{a_{d}\cdot{\left|\mathsf{X}[k]\right|}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]\to\sqrt{2}\sigma,$ (E.46)
which proves (E.8).
Proof of (E.26) and (E.9).
The numerator in (C.8) converges to,
$\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]}^{2}=\frac{1}{2}\left[\mathbb{E}\left|\mathsf{N}[k]\right|^{2}-\mathbb{E}[\left|\mathsf{N}[k]\right|^{2}\cos(2\phi_{e}[k])]\right]\rightarrow\frac{1}{2}\mathbb{E}\left|\mathsf{N}[k]\right|^{2},$ (E.47)
$\frac{1}{\sigma^{2}}\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]}^{2}\rightarrow\frac{1}{2\sigma^{2}}\mathbb{E}\left|\mathsf{N}[k]\right|^{2}=1,$ (E.48)
which concludes the proof.
Appendix F Proof of Theorem 4.3
To begin, let us summarize the notation and results from the previous sections as the foundation for the proof. Recall the definition of $\phi_{e}[k]$ in (A.7), as well as the definitions of $f_{1}\left(r\right)$ and $f_{2}\left(r\right)$ in (E.6)-(E.7), and let $a_{d}\triangleq\sqrt{2\log d}$. According to Theorem 4.1, the convergence of the Fourier phases is given by:
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=C_{k}.$ (F.1)
where $C_{k}$ is given by (C.8):
$C_{k}=\frac{{\mathbb{E}\left[\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]^{2}\right]}}{{\left(\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\cos(\phi_{e}[k])\right]}\right)}^{2}}.$ (F.2)
The constant $C_{k}$ can be rewritten as:
$a^{2}_{d}\cdot C_{k}=a_{d}^{2}\cdot\frac{{\mathbb{E}\left[\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]^{2}\right]}}{{\left(\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\cos(\phi_{e}[k])\right]}\right)}^{2}}=a_{d}^{2}\cdot\frac{\mathbb{E}{[{f_{2}(\hat{\mathsf{R}})}]}}{\left(\mathbb{E}{[{f_{1}(\hat{\mathsf{R}})}]}\right)^{2}},$ (F.3)
where $f_{1}\left(r\right)$ and $f_{2}\left(r\right)$ are defined in (E.6)-(E.7).
The signal $x$, satisfying the conditions of Theorem 4.3, also satisfies the assumptions of Lemma E.1. By Lemma E.1, we have:
$\lim_{d\to\infty}\ \frac{1}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\sqrt{2}\sigma,$ (F.4)
and,
$\lim_{d\to\infty}\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\sigma^{2}.$ (F.5)
@@
From (F.3) and (F.5), we obtain:
$\lim_{d\to\infty}a^{2}_{d}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}=\frac{1}{2}.$ (F.6)
Combining (F.1), (F.6), results,
$\lim_{d\to\infty}\lim_{M\to\infty}$ $\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}\frac{1}{1/(4\log\left(d\right)\left|\mathsf{X}[k]\right|^{2})}=\lim_{d\to\infty}\frac{C_{k}}{1/(4\log\left(d\right)\left|\mathsf{X}[k]\right|^{2})}$ (F.7)
$=\lim_{d\to\infty}\frac{a_{d}^{2}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}}{1/2}=1,$ (F.8)
where (F.7) follows from (F.1), and (F.8) follows from (F.6), proving (4.6).
@@
$|\mathsf{\hat{X}}[k]|\xrightarrow[]{\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]=\mathbb{E}{[{f_{1}(\hat{\mathsf{R}})}]}.$ (F.9)
Combining (F.4), (F.9) yields,
$\frac{1}{\sqrt{2}a_{d}\sigma}\frac{|\mathsf{\hat{X}}[k]|}{\left|\mathsf{X}[k]\right|}\xrightarrow[]{\mathsf{a.s.}}\frac{1}{\sqrt{2}a_{d}\sigma}\frac{\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]}{\left|\mathsf{X}[k]\right|}\rightarrow 1,$ (F.10)
as $M,d\to\infty$, where the second passage follows from (E.8). As $a_{d}=\sqrt{2\log\left(d\right)}$, this completes the proof of the Theorem.
Appendix G Proof of Proposition 5.1
@@
$\mathbb{E}\left[\max\{A_{0},A_{1},\ldots,A_{d-1}\}\right]\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}\left[|A_{r_{1}}-A_{r_{2}}|\right].$ (G.1)
Proof of Lemma G.1.
@@
$\mathbb{E}[\max\{A_{r_{1}},A_{r_{2}}\}]$ $=\frac{1}{2}\mathbb{E}[A_{r_{1}}+A_{r_{2}}+|A_{r_{1}}-A_{r_{2}}|]$ (G.3)
$=\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|],$ (G.4)
@@
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]\geq\mathbb{E}[\max\{A_{r_{1}},A_{r_{2}}\}],$ (G.5)
for every $r_{1},r_{2}\in\{0,1,\ldots,d-1\}$. Combining (G.4) and (G.5), we conclude:
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|],$ (G.6)
@@
$\hat{x}\triangleq\frac{1}{M}\sum_{i=0}^{M-1}\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},$ (G.7)
@@
$\mathsf{\hat{R}}_{i}\triangleq\operatorname*{arg\,max}_{0\leq\ell\leq d-1}\langle n_{i},\mathcal{T}_{\ell}x\rangle.$ (G.8)
@@
$\langle\hat{x},x\rangle$ $=\left\langle\frac{1}{M}\sum_{i=0}^{M-1}\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},x\right\rangle=\frac{1}{M}\sum_{i=0}^{M-1}\langle\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},x\rangle.$ (G.9)
@@
[…]
@@
$A_{r}\triangleq\langle n_{1},\mathcal{T}_{r}x\rangle.$ (G.11)
Then, the right-hand side of (G.10) becomes:
$\mathbb{E}[\langle\mathcal{T}_{-\mathsf{\hat{R}}_{1}}n_{1},x\rangle]=\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}].$ (G.12)
Applying Lemma G.1, we get:
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]$ $\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|]$ (G.13)
$=\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|\langle n_{1},\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\rangle|].$ (G.14)
To complete the proof, we show that the lower bound in (G.14) is strictly positive. Since $x\in\mathbb{R}^{d}$ is nonzero with non-vanishing Fourier components $\mathsf{X}[k]\neq 0$, for every $11\leq k\leq d-$ 1, and $\mathcal{T}_{r}$ is a cyclic shift operator, the set $\{\mathcal{T}_{r}x:0\leq r<d\}$ contains at least $d-1$ distinct vectors. Thus, there exist $r_{1},r_{2}\in\{0,\ldots,d-1\}$ such that
$v\triangleq\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\neq 0.$ (G.15)
@@
$\mathrm{Var}(\langle n_{1},v\rangle)=\mathbb{E}[\langle n_{1},v\rangle^{2}]=v^{\top}\Sigma v>0,$ (G.16)
@@
$\mathbb{E}[|\langle n_{1},v\rangle|]>0.$ (G.17)
@@
$\max_{0\leq r_{1},r_{2}<d}\frac{1}{2}\mathbb{E}[|\langle n_{1},\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\rangle|]>0,$ (G.18)
@@
$\lim_{M\to\infty}\langle\hat{x},x\rangle=\mathbb{E}[\langle\mathcal{T}_{-\hat{\mathsf{R}}_{1}}n_{1},x\rangle]>0,$ (G.19)
@@
Appendix H Proof of Theorem 5.2
In this section, we prove Theorem 5.2. The proof relies on the functional central limit theorem for the discrete Fourier transform [34, 12, 11], which we review in Appendix H.1. In Appendices H.2 and H.3, we apply this result to analyze the real and imaginary parts of the EfN estimator under a general i.i.d. noise model, and compare the outcome to the white Gaussian case. Finally, the proof of Theorem 5.2 is deduced in Appendix H.4.
H.1 The functional CLT for DFT

### v3

@@ D.2 High-dimensional limits
$\frac{\sigma_{k}^{2}}{2\left|\mathsf{X}[k]\right|^{2}a_{d}}\mathbb{E}\left[\left.{\widetilde{\mu}_{k}[\hat{\mathsf{R}}]}\right|\mathsf{N}[k]\right]\xrightarrow[]{\mathsf{a.s.}}\left|\mathsf{N}[k]\right|^{2}.$ (D.42)
Let the term on the left-hand side of (D.42) be denoted by $G_{d}$ and the term on the right-hand side by $G$, so that $G_{d}\xrightarrow[]{\text{a.s.}}G$. By definition, observe that $\left|G_{d}\right|\leq\left|\mathsf{N}[k]\right|^{2}$, and it is evident that $\mathbb{E}[\left|\mathsf{N}[k]\right|^{2}]=\sigma^{2}<\infty$. Thus, by the dominated convergence theorem, we conclude that $G_{d}\xrightarrow[]{\mathcal{L}^{1}}G$, and specifically,
$\mathbb{E}\left[\frac{\sigma_{k}^{2}}{2a_{d}}\frac{\mathbb{E}[\mu(\hat{\mathsf{R}})|\mathsf{N}[k]]}{\left|\mathsf{X}[k]\right|^{2}}\right]\rightarrow\mathbb{E}\left[\left|\mathsf{N}[k]\right|^{2}\right].$ (D.43)
By combining (D.43) with (D.29), we get,
$\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}{\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]}\xrightarrow[]{\mathcal{L}^{1}}\left|\mathsf{N}[k]\right|^{2},$ (D.44)
or equivalently,
$\mathbb{E}\left[\frac{\sigma_{k}}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}\left[\left.{f_{1}(\hat{\mathsf{R}})}\right|\mathsf{N}[k]\right]\right]\to\mathbb{E}\left[\left|\mathsf{N}[k]\right|^{2}\right],$ (D.45)
which proves (D.26).
By the law of total expectation, we then obtain
$\frac{\sigma_{k}}{a_{d}\cdot{\left|\mathsf{X}[k]\right|}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]\to\mathbb{E}\left[{\left|\mathsf{N}[k]\right|^{2}}\right]=\sigma^{2},$ (D.46)
as $d\to\infty$. By Lemma D.1, we know that $\sigma_{k}^{2}\to\sigma^{2}$ as $d\to\infty$, so from (D.46), we conclude that
$\frac{1}{a_{d}\cdot{\left|\mathsf{X}[k]\right|}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]\to\sigma,$ (D.47)
which proves (D.22).
Proof of (D.27) and (D.23).
The numerator in (B.64) converges to,
$\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]}^{2}=\frac{1}{2}\left[\mathbb{E}\left|\mathsf{N}[k]\right|^{2}-\mathbb{E}[\left|\mathsf{N}[k]\right|^{2}\cos(2\phi_{e}[k])]\right]\rightarrow\frac{1}{2}\mathbb{E}\left|\mathsf{N}[k]\right|^{2},$ (D.48)
$\frac{1}{\sigma^{2}}\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]}^{2}\rightarrow\frac{1}{2\sigma^{2}}\mathbb{E}\left|\mathsf{N}[k]\right|^{2}=\frac{1}{2},$ (D.49)
which concludes the proof. ∎
D.3 Proof of Theorem 4.3
To begin, let us summarize the notation and results from the previous sections as the foundation for the proof. Recall the definition of $\phi_{e}[k]$ in (A.7), as well as the definitions of $f_{1}\left(r\right)$ and $f_{2}\left(r\right)$ in (D.20)-(D.21), and let $a_{d}\triangleq\sqrt{2\log d}$. According to Theorem 4.1, the convergence of the Fourier phases is given by:
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=C_{k}.$ (D.50)
where $C_{k}$ is given by (B.64):
$C_{k}=\frac{{\mathbb{E}\left[\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]^{2}\right]}}{{\left(\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\cos(\phi_{e}[k])\right]}\right)}^{2}}.$ (D.51)
The constant $C_{k}$ can be rewritten as:
$a^{2}_{d}\cdot C_{k}=a_{d}^{2}\cdot\frac{{\mathbb{E}\left[\left[\left|\mathsf{N}[k]\right|\sin(\phi_{e}[k])\right]^{2}\right]}}{{\left(\mathbb{E}{\left[\left|\mathsf{N}[k]\right|\cos(\phi_{e}[k])\right]}\right)}^{2}}=a_{d}^{2}\cdot\frac{\mathbb{E}{[{f_{2}(\hat{\mathsf{R}})}]}}{\left(\mathbb{E}{[{f_{1}(\hat{\mathsf{R}})}]}\right)^{2}},$ (D.52)
where $f_{1}\left(r\right)$ and $f_{2}\left(r\right)$ are defined in (D.20)-(D.21).
The signal $x$, satisfying the conditions of Theorem 4.3, also satisfies the assumptions of Proposition D.3. By Proposition D.3, we have:
$\lim_{d\to\infty}\ \frac{1}{a_{d}\cdot\left|\mathsf{X}[k]\right|}\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]=\sigma,$ (D.53)
@@
$\lim_{d\to\infty}\mathbb{E}[{f_{2}(\hat{\mathsf{R}})}]=\frac{\sigma^{2}}{2}.$ (D.54)
@@
From (D.52)–(D.54), we obtain:
$\lim_{d\to\infty}a^{2}_{d}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}=\frac{1}{2}.$ (D.55)
Combining (D.50), (D.55), results,
$\lim_{d\to\infty}\lim_{M\to\infty}$ $\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}\frac{1}{1/(4\log\left(d\right)\left|\mathsf{X}[k]\right|^{2})}=\lim_{d\to\infty}\frac{C_{k}}{1/(4\log\left(d\right)\left|\mathsf{X}[k]\right|^{2})}$ (D.56)
$=\lim_{d\to\infty}\frac{a_{d}^{2}\cdot\left|\mathsf{X}[k]\right|^{2}\cdot C_{k}}{1/2}=1,$ (D.57)
where (D.56) follows from (D.50), and (D.57) follows from (D.55), proving (4.8).
@@
$|\mathsf{\hat{X}}[k]|\xrightarrow[]{\mathsf{a.s.}}\mathbb{E}\left[\left|\mathsf{N}[k]\right|\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}+\phi_{\mathsf{N}}[k]-\phi_{\mathsf{X}}[k]\right)\right]=\mathbb{E}{[{f_{1}(\hat{\mathsf{R}})}]}$. (D.58)
Combining (D.53), (D.58) yields,
$\frac{1}{a_{d}\sigma}\frac{|\mathsf{\hat{X}}[k]|}{\left|\mathsf{X}[k]\right|}\xrightarrow[]{\mathsf{a.s.}}\frac{1}{a_{d}\sigma}\frac{\mathbb{E}[{f_{1}(\hat{\mathsf{R}})}]}{\left|\mathsf{X}[k]\right|}\rightarrow 1,$ (D.59)
as $M,d\to\infty$, where the second passage follows from (D.22). As $a_{d}=\sqrt{2\log\left(d\right)}$, this completes the proof of the Theorem.
Appendix E Proof of Proposition 5.1
@@
$\mathbb{E}\left[\max\{A_{0},A_{1},\ldots,A_{d-1}\}\right]\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}\left[|A_{r_{1}}-A_{r_{2}}|\right].$ (E.1)
Proof of Lemma E.1.
@@
$\mathbb{E}[\max\{A_{r_{1}},A_{r_{2}}\}]$ $=\frac{1}{2}\mathbb{E}[A_{r_{1}}+A_{r_{2}}+|A_{r_{1}}-A_{r_{2}}|]$ (E.3)
$=\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|],$ (E.4)
@@
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]\geq\mathbb{E}[\max\{A_{r_{1}},A_{r_{2}}\}],$ (E.5)
for every $r_{1},r_{2}\in\{0,1,\ldots,d-1\}$. Combining (E.4) and (E.5), we conclude:
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|],$ (E.6)
@@
$\hat{x}\triangleq\frac{1}{M}\sum_{i=0}^{M-1}\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},$ (E.7)
@@
$\mathsf{\hat{R}}_{i}\triangleq\operatorname*{arg\,max}_{0\leq\ell\leq d-1}\langle n_{i},\mathcal{T}_{\ell}x\rangle.$ (E.8)
@@
$\langle\hat{x},x\rangle$ $=\left\langle\frac{1}{M}\sum_{i=0}^{M-1}\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},x\right\rangle=\frac{1}{M}\sum_{i=0}^{M-1}\langle\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},x\rangle.$ (E.9)
@@
$\frac{1}{M}\sum_{i=0}^{M-1}\langle\mathcal{T}_{-\mathsf{\hat{R}}_{i}}n_{i},x\rangle\xrightarrow[]{\text{a.s.}}\mathbb{E}[\langle\mathcal{T}_{-\mathsf{\hat{R}}_{1}}n_{1},x\rangle].$ (E.10)
@@
$A_{r}\triangleq\langle n_{1},\mathcal{T}_{r}x\rangle.$ (E.11)
Then, the right-hand side of (E.10) becomes:
$\mathbb{E}[\langle\mathcal{T}_{-\mathsf{\hat{R}}_{1}}n_{1},x\rangle]=\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}].$ (E.12)
Applying Lemma E.1, we get:
$\mathbb{E}[\max\{A_{0},A_{1},\ldots,A_{d-1}\}]$ $\geq\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|A_{r_{1}}-A_{r_{2}}|]$ (E.13)
$=\max_{0\leq r_{1},r_{2}\leq d-1}\frac{1}{2}\mathbb{E}[|\langle n_{1},\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\rangle|].$ (E.14)
To complete the proof, we show that the lower bound in (E.14) is strictly positive. Since $x\in\mathbb{R}^{d}$ is nonzero with non-vanishing Fourier components $\mathsf{X}[k]\neq 0$, for every $1\leq k\leq d-1$, and $\mathcal{T}_{r}$ is a cyclic shift operator, the set $\{\mathcal{T}_{r}x:0\leq r<d\}$ contains at least d1d-1 distinct vectors. Thus, there exist $r_{1},r_{2}\in\{0,\ldots,d-1\}$ such that
$v\triangleq\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\neq 0.$ (E.15)
@@
$\mathrm{Var}(\langle n_{1},v\rangle)=\mathbb{E}[\langle n_{1},v\rangle^{2}]=v^{\top}\Sigma v>0,$ (E.16)
@@
$\mathbb{E}[|\langle n_{1},v\rangle|]>0.$ (E.17)
@@
$\max_{0\leq r_{1},r_{2}<d}\frac{1}{2}\mathbb{E}[|\langle n_{1},\mathcal{T}_{r_{1}}x-\mathcal{T}_{r_{2}}x\rangle|]>0,$ (E.18)
@@
$\lim_{M\to\infty}\langle\hat{x},x\rangle=\mathbb{E}[\langle\mathcal{T}_{-\hat{\mathsf{R}}_{1}}n_{1},x\rangle]>0,$ (E.19)
@@
Appendix F Proof of Theorem 5.2: High-dimensional i.i.d. noise
In this section, we prove Theorem 5.2. The proof relies on the functional central limit theorem for the discrete Fourier transform [37, 14, 13], which we review in Appendix F.1. In Appendices F.2 and F.3, we apply this result to analyze the real and imaginary parts of the EfN estimator under a general i.i.d. noise model, and compare the outcome to the white Gaussian case. Finally, the proof of Theorem 5.2 is deduced in Appendix F.4.
F.1 The functional CLT for DFT